OpenStudy (anonymous):

for interval [0,2π) solve cos(4x)=1/2

OpenStudy (anonymous):

4x =arccos(1/2)

OpenStudy (anonymous):

x= pi/12

OpenStudy (anonymous):

+πk/2 k=0,1,2,3?

OpenStudy (anonymous):

wait we need to continue

OpenStudy (anonymous):

(0,2pi), now we know that cos is positive in in the first and forth quarter

OpenStudy (anonymous):

First solve \(\cos(\theta)=1/2, \theta\in [0,2\pi)\)

OpenStudy (anonymous):

This is an easy angle.  Actually \(\theta \in [0,8\pi)\) So we have \(\theta = \pi/3,\theta = 2\pi - \pi/3,...\)

OpenStudy (anonymous):

\[ \theta = \{\pi/3,5\pi/3,7\pi/3,11\pi/3,13\pi/3,17\pi/3,19\pi/3,23\pi/3\} \]Now \[ 4x = \theta \implies x = \theta/4 \]So divide all those solutions by 4